# Consistent Validation for Predictive Methods in Spatial Settings

### v1

@@ 4 Current Estimators Exhibit Inconsistency
Even though consistency under infill asymptotics is a minimal property, we next prove that principle existing test-risk estimators fail to satisfy it.
4.1 Inconsistency of the Holdout
The holdout estimator of test risk is well-justified and consistent in the traditional case of i.i.d. test and validation data (Devroye, 1976; Langford, 2005). But we next show that it need not be consistent for spatial problems.
Proposition 4.1 (Counterexample to the consistency of holdout).
Consider a spatial location $S$ and a single test point at this location $(S,X,Y)$, with covariate $X$ as at the start of Section 2 and response $Y$ as in Assumption 2.2. Suppose the test risk at such a point (Def. 2.3 with $M^{\textup{{test}}}=1$) is not constant as a function of $S$. Then there exists a test set (of size one) such that $\hat{R}^{\chi}_{\textup{{Hold}}}$ is not a consistent estimator of the test risk on that test set.
Intuition for Inconsistency of the Holdout.
Prop. 4.1 relies on the key observation that $\hat{R}^{\chi}_{\textup{{Hold}}}$ has no dependence on the test task. So unless all test tasks have the same risk (which will be true only in unusually simplistic settings), it cannot estimate them all consistently. One might think of the holdout estimator as generally exhibiting non-trivial bias since it averages loss across the validation sites when we really care about loss at the test sites. We give a formal proof following this intuition in Section B.3.1.
4.2 Inconsistency of 1-Nearest Neighbor
Because of the assumed regularity in the error function (Assumption 2.5), it is natural to estimate the error at a test site using nearby validation points. Loog (2012) proposed risk estimators using $k$-nearest neighbors in the context of covariate shift. Both Loog (2012) and Portier et al. (2023) advocated for the use of 1-nearest neighbor (1NN) in the covariate-shift setting, with the latter providing theoretical justifications under standard covariate-shift assumptions. However, we next produce common spatial examples where the 1NN estimator exhibits inconsistency.
Nearest Neighbor Estimator.
We start by reviewing a general $k$-nearest neighbor estimator, which we revisit later. Define the $k$-nearest neighbor radius of a point $S\in\mathcal{S}$ as $\tau^{k}(S)=\inf\{a\in\mathbb{R}:|S^{\textup{{val}}}_{1:N^{\textup{{val}}}}\cap B(S,a)|\geq k\},$ where $B(S,a)$ denotes the ball of radius $a$ centered at $S$. Then the $k$-nearest neighbor set [^5] We will state our results for the version of nearest neighbors where ties are resolved by including all equidistant points. However, our analysis holds for arbitrary tie-breaking methods. of a point $S\in\mathcal{S}$ is
$A^{k}(S)=\{1\leq n\leq N^{\textup{{val}}}:\,S^{\textup{{val}}}_{n}\in B(S,\tau^{k}(S))\}.\,$ (6)
As long as $1\leq k\leq N^{\textup{{val}}}$, $A^{k}(S)$ contains at least $k$ points: the $k$ nearest neighbors to $S$ in the validation set. It may be larger than $k$ if multiple points are equidistant from $S$.
@@ 4.2 Inconsistency of 1-Nearest Neighbor
The $k$-nearest neighbor test-risk estimator is defined by
$\hat{R}^{\chi}_{w^{\textup{{NN}},k}}$ $(h)=\sum\nolimits_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n})),\textrm{ where}$ (7)
$\!w^{\textup{{NN}},k}_{n}$ $\!=\!\tfrac{1}{M^{\textup{{test}}}}\sum\nolimits_{m=1}^{M^{\textup{{test}}}}\tfrac{1}{|A^{k}(S^{\textup{{test}}}_{m})|}\mathrm{1}\{S^{\textup{{val}}}_{n}\!\in\!A^{k}(S^{\textup{{test}}}_{m})\}.$ (8)
@@
Loog (2012) proposed weighting the loss function in this way when training a model under covariate shift. Portier et al. (2023) analyzed a similar approach, in which validation points are sampled with probabilities corresponding to the weights in Eqn. 8, for mean estimation. Portier et al. (2023) made standard covariate shift assumptions of i.i.d. validation sites, i.i.d. test sites, and a bounded density ratio between the validation and test distributions.
Our inconsistency result (Prop. 4.3 below) relies on the observation that, unlike in the covariate-shift setting, it is often the case in the spatial setting that test points are fixed rather than arising i.i.d. from a distribution. We first state our result and then discuss further intuition.
Proposition 4.3 (Counterexample to the consistency of 1NN).
For a particular spatial location $S$, let the associated test point be $(S,X,Y)$, with covariate $X$ as at the start of Section 2 and response $Y$ as in Assumption 2.2. Assume there exists a constant $c>0$ such that for all $S\in\mathcal{S}$ and associated $(S,X,Y)$, $\mathrm{Var}[\ell(Y,h^{\chi}(S))|S]\geq c$. Next, consider an infinite sequence of validation sets as in Def. 3.2. Suppose there exists an $S^{\prime}\in\mathcal{S}$ such that for any $r>0$ and $N^{\textup{{val}}}>0$, |$|\{1\leq j\leq N^{\textup{{val}}}:d_{\mathcal{S}}(S^{\textup{{val}}}_{j},S^{\prime})=r\}|\leq 1$. Choose any $Q^{\textup{{test}}}$ such that $Q^{\textup{{test}}}(\{S^{\prime}\})>0$. Then there exists a $\delta\in(3/4,1)$ and a $C^{(\delta)}>0$ such that, for each $N^{\textup{{val}}}$, with probability at least $1-\delta$, $|R^{\chi}_{Q^{\textup{{test}}}}(h)\!-\!(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}(h)|\geq C^{(\delta)}.$ Here $(\hat{R}^{\chi}_{w^{\textup{{NN}},\!1}})^{(N^{\textup{{val}}})}$ denotes the 1NN estimator associated to the first $N^{\textup{{val}}}$ data points and $\delta$. $C^{(\delta)}$ and $\delta$ do not depend on $N^{\textup{{val}}}$ or other properties of the sequence of validation data.
See Section B.3.2 for a proof of the result. The technical condition |$|\{1\leq n\leq N^{\textup{{val}}}:d_{\mathcal{S}}(S^{\textup{{val}}}_{j},S^{\prime})=r\}|\leq 1$ ensures that the 1NN set for each point contains exactly $1$ point. That is, there are no ties. This condition would be satisfied for the infinite sequence of validation sets with probability one if, for instance, the validation points were chosen i.i.d. from a uniform measure on a compact set; cf. Prop. 3.3.
Intuition for Inconsistency of $1$-Nearest Neighbor. The variance of the 1NN estimator will, in non-trivial cases, remain non-negligible even as the amount of validation data increases without bound. Consider the simple case where $Q^{\textup{{test}}}=\delta_{S}$ for some $S\in\mathcal{S}$. Using $k=1$ would lead us to estimate the error using a single validation point, which is not consistent due to the noise in the observed response at the validation point. Where the problem with the holdout estimator was bias, the problem with 1NN is variance.
@@ 5 A Consistent Estimator
We next provide a novel bound on the test-risk estimation error of $k$-nearest neighbors. We propose using a $k$-nearest neighbors estimator with $k$ chosen by optimizing our bound. We show that our proposed estimator is consistent for test risk under infill asymptotics. Finally, we provide promising results for model selection but also discuss open challenges.
@@ 5.1 Our Bound and Estimator
Analogous to the intuition after Prop. 4.3, we expect inconsistency of $k$-nearest neighbors for any fixed choice of $k$. Instead, we propose to trade off the larger variance of small $k$ and larger bias of large $k$ by optimizing a bound depending on the validation set. To that end, we first derive a bound on the test-risk estimation error as a function of $k$.
@@
$\zeta^{k}(\Psi_{1};\Psi_{2})=\sup_{S_{2}\in\Psi_{2}}\,\inf_{\begin{subarray}{c}A\subset\Psi_{1},\\ |A|=k\end{subarray}}\,\,\sup_{S_{1}\in A}d_{\mathcal{S}}(\Psi_{1},\Psi_{2}).$ (11)
@@
Consider a validation set $D^{\textup{{val}}}$ of size $N^{\textup{{val}}}$ and a test set $D^{\textup{{test}}}$ of size $M^{\textup{{test}}}$. Take the $k$-nearest neighbors test-risk estimator from Def. 4.2. Choose $k$ such that $1\leq k\leq N^{\textup{{val}}}$, $\delta\in(0,1)$, and $\rho_{k}\coloneqq\zeta^{k}(S^{\textup{{val}}}_{1:N^{\textup{{val}}}},S^{\textup{{test}}}_{1:M^{\textup{{test}}}})$. Take Assumptions 2.2, 2.1, 2.4 and 2.5. Then, with probability at least $1-\delta$,
[…]
$\leq L\rho_{k}\!+\!\Delta\sqrt{\!\max_{1\leq n\leq N^{\textup{{val}}}}\frac{Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))}{k}}\sqrt{\tfrac{1}{2}\log\tfrac{2}{\delta}},$ (13)
@@
where $B(S,r)$ denotes the ball of radius $r$ centered at $S$. See Assumptions 2.1, 2.5 and 1 for $\Delta$,$L$, $Q^{\textup{{test}}}$ respectively.
@@
We prove Thm. 5.1 in Section B.4. Roughly the first term on the righthand side in Eqns. 12 and 13 is a bias term, and the second term is a variance term. Specifically, the first term is a worst-case upper bound on the bias of our estimator; it will be large if the average loss varies quickly in space or if validation data is not available near test data. Larger $k$ may increase the first term. The second term comes from applying a tail bound; if most of the weight is put on a few validation points, the resulting estimator has high variance and this term is large. Sufficiently large $k$ will decrease the second term. If we can find a $k$ such that both (a) the distance from each test point to its $k$ nearest neighbors is small and (b) no validation point has too much impact on our estimator, $\hat{R}^{\chi}_{w^{\textup{{NN}},k}}(h)$ provides a good estimate for $R^{\chi}_{Q^{\textup{{test}}}}$.
@@
Thm. 5.1 is closely related to Portier et al. (2023, Prop. 4). While there are technical differences in the proof and algorithm (and the risk that is bounded), the substantive distinction is that we state our bound directly in terms of the fill distance, instead of upper bounding this distance again as done in Portier et al. (2023). We can therefore avoid making assumptions about the distributions of the sites; we instead highlight the fill distance of the validation set as an essential quantity in controlling the accuracy of nearest neighbor risk estimation.
@@
Selection Procedure with Unknown Lipschitz Constant. If the Lipschitz constant of the average loss, $L$, can be upper bounded, for example by knowledge about how quickly varying the spatial processes involved in the analysis are, then $k$ can be selected by minimizing the upper bound in Eqn. 12. Since the bound is conditional on the validation and test sites, the bound still holds with the same probability for $k$ selected by this minimization. However, it will generally be the case that the Lipschitz constant is unknown. We therefore suggest choosing the number of neighbors by minimizing the upper bound from Thm. 5.1 with $1$ in place of the Lipschitz constant:
$k^{\star}_{T}\in\arg\min\nolimits_{k\in T}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2},$ (14)
@@
with $\beta_{\delta}=\smash{\Delta\sqrt{\frac{1}{2}\log\frac{2}{\delta}}}$. For computational efficiency, we focus on choosing $k$ as a power of 2: $T=T_{2}:=\{2^{i}\}_{i=1}^{\smash{\lfloor\log_{2}N^{\textup{{val}}}\rfloor}}$. We call the resulting estimator spatial nearest neighbors (SNN).
@@ 5.2 Our Method is Consistent
Suppose that $\mathcal{S}=[0,1]^{d}$. Take Assumptions 2.2, 2.1, 2.4 and 2.5. Let $k^{\star}_{T_{2}}\in\arg\min_{k\in T_{2}}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}$ with $\beta_{\delta}=\Delta\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$. Then $\hat{R}^{\chi}_{w^{\textup{{{NN}}},k^{\star}_{T_{2}}}}(h)$ is consistent under infill asymptotics.
See Section B.5 for a proof of Cor. 5.2. Cor. 5.2 tells us that minimizing an upper bound on our error in estimation leads to an estimator that is consistent regardless of the test data, as long as the validation data are dense on the unit cube.
5.3 Model Selection
(1) Fixed predictive method. So far we have focused on the setting where predictive methods are fixed in advance. As we accrue validation data, the consistency of SNN ensures it will eventually choose the predictive method with lowest test risk. Without consistency, holdout and 1NN cannot be trusted to choose the best predictive method.
We are not able to formally characterize how SNN performs in model selection. But we provide rigorous results on rates of convergence of SNN that give suggestive guidance. In particular, (a) we first consider the case of a grid prediction task, where test sites are arranged on a grid. In this case, we are able to give a rate of convergence for both $1$NN and SNN (Cors. B.12 and B.13). If the validation data are i.i.d. our rate is faster than the minimax optimal convergence rate for predicting a Lipschitz function in the presence of additive, homoskedastic noise. Our result shows the same rate of convergence as Portier et al. (2023, Prop. 4), who considered test data that was i.i.d. instead of on a grid. Our result is suggestive that both nearest neighbor methods may perform well at model selection for validating maps.
(b) Second, we provide finite-sample bounds and asymptotic characterizations for general (non-grid) $Q^{\textup{{test}}}$ (Cor. B.10). In this case, when validation data is i.i.d., up to logarithmic factors, we show that SNN converges at the optimal rate of convergence for Lipschitz functions (Cor. B.14). Since in this case, the SNN convergence rate is not strictly faster than the training convergence rate, it might be difficult to select between statistical methods that converge at the optimal regression rate. But we would still expect to be able to select between (i) a statistical method that converges at the optimal rate and (ii) one that does not (for example a misspecified parametric model). Since the holdout and 1NN may not even be consistent, we could not rely on them to select the better model even in this latter, easier case.

### v2

@@ 4 Current Estimators Exhibit Inconsistency
Even though consistency under infill asymptotics is a minimal desirable property, we next prove that principle existing test-risk estimators fail to satisfy it in realistic problems.
Inconsistency of the Holdout. We state our result and then discuss the realism of the example.
Proposition 4.1 (Inconsistency of holdout).
There exists a set of test points and a data-generating process satisfying infill asymptotics such that $\smash{\hat{R}_{\textup{{Hold}}}}$ is not a consistent estimator of the test risk.
While the holdout estimator of test risk is consistent for i.i.d. test and validation data (Devroye, 1976; Langford, 2005), we can construct examples showing Proposition 4.1 by observing that $\smash{\hat{R}_{\textup{{Hold}}}}$ has no dependence on the test task. So unless all test tasks have the same risk (which will be true only in unusually simplistic spatial settings), it cannot estimate them all consistently. The holdout estimator will generally exhibit non-trivial bias since it averages loss across the validation sites when we really care about loss at the test sites. See Section D.3.1 for a formal proof and also an example where the holdout converges to $\Delta$, the maximum possible error under the loss bound.
Nearest Neighbor Estimator. Because of the assumed regularity in the error function (Assumption 2.5), it is natural to estimate the error at a test site using nearby validation points. Loog (2012) proposed risk estimators using $k$-nearest neighbors in the context of covariate shift. Both Loog (2012) and Portier et al. (2023) advocated for the use of 1-nearest neighbor (1NN) in the covariate-shift setting, with the latter providing theoretical justifications under standard covariate-shift assumptions. However, we show the 1NN estimator exhibits inconsistency in our spatial setting.
We first review a general $k$-nearest neighbor estimator, which we revisit later. Define the $k$-nearest neighbor radius of a point $S\!\in\!\mathcal{S}$ as $\!\tau^{k}(S)\!:=\!\inf\{a\!\in\!\mathbb{R}\!:\!\!|S^{\textup{{val}}}_{1:N^{\textup{{val}}}}\!\cap\!B(S,a)|\!\geq\!k\},$ where $B(S,a)$ is the ball of radius $a$ centered at $S$. The $k$-nearest neighbor set [^4] We will state our results for the version of nearest neighbors where ties are resolved by including all equidistant points. However, our analysis holds for arbitrary tie-breaking methods. of a point $S\in\mathcal{S}$ is $A^{k}(S)\!:=\!\{1\!\leq n\!\leq\!N^{\textup{{val}}}\!:\!S^{\textup{{val}}}_{n}\in\!B(S,\!\tau^{k}(S))\}.$ As long as $1\leq k\leq N^{\textup{{val}}}$, $A^{k}(S)$ contains at least $k$ points: the $k$ nearest neighbors to $S$ in the validation set. It may be larger than $k$ if multiple points are equidistant from $S$.
@@
The $k$-nearest neighbor (kNN) test-risk estimator is defined by $\smash{\hat{R}_{\textup{{NN}},k}}(h):=\sum\nolimits_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}_{n}\ell(Y^{\textup{{val}}}_{n},h^{\chi}(S^{\textup{{val}}}_{n}))$, where $w^{\textup{{NN}},k}_{n}:=(1/M^{\textup{{test}}})\sum\nolimits_{m=1}^{M^{\textup{{test}}}}\mathrm{1}\{S^{\textup{{val}}}_{n}\!\in\!A^{k}(S^{\textup{{test}}}_{m})\}/|A^{k}(S^{\textup{{test}}}_{m})|.$
@@
Loog (2012) proposed weighting the loss function in this way when training a model under covariate shift. Portier et al. (2023) analyzed a similar approach, in which validation points are sampled with probabilities corresponding to the weights in Definition 4.2, for mean estimation. Portier et al. (2023) made standard covariate shift assumptions of i.i.d. validation sites, i.i.d. test sites, and a bounded density ratio between the validation and test distributions.
Inconsistency of 1NN. We again state our result and then develop intuition.
Proposition 4.3 (Inconsistency of 1NN).
There exist a set of test points and a data-generating process satisfying infill asymptotics such that $\hat{R}_{\textup{{NN}},1}(h)$ is not a consistent estimator of the test risk.
For intuition, recall that—unlike in the covariate-shift setting—test points in the spatial setting are commonly fixed rather than arising i.i.d. from a distribution. Consider the simple case where $Q^{\textup{{test}}}=\delta_{S}$ for some $S\in\mathcal{S}$. Using $k=1$ leads us to estimate the error using a single validation point, which is inconsistent due to observation noise at the validation point. Where the problem with the holdout estimator was bias, the problem with 1NN is variance. In Section D.3.2 we prove Proposition 4.3 and show 1NN has large error when applied to classification point prediction tasks.
Inconsistency of Nearest Neighbors When $k$ Is a Function of the Number of Validation Points. In fact, we can prove a more general result: that any nearest-neighbor test-risk estimator where the number of neighbors depends (only) on the number of validation points is inconsistent under infill asymptotics, regardless of type of dependence.
Proposition 4.4 (Inconsistency of kNN depending on number of validation points).
Let $(k_{n})_{n=1}^{\infty}$ be any sequence of natural numbers. Define the sequence of estimators $\smash{\hat{R}_{N^{\textup{{val}}}}}$ to be the nearest neighbor risk estimators using $N^{\textup{{val}}}$ validation points and $k_{N^{\textup{{val}}}}$ neighbors. Then there exists a data-generating process satisfying infill asymptotics, a test set containing a single point, a predictive method $h$ resulting in an error function satisfying the Lipschitz assumption, and an $\!\epsilon,\delta>\!0$ such that with probability at least $1\!-\!\delta$, $\forall N^{\textup{{val}}}$, $|\smash{\hat{R}_{N^{\textup{{val}}}}}(h)-R_{Q^{\textup{{test}}}}(h)|\geq\epsilon$.
See Section D.3.5 for a proof. There are two cases. (1) If the number of neighbors is bounded, the estimator suffers from non-vanishing variance as in the 1NN case. Or (2) the number is unbounded, so there exists a sequence of validation sites that accumulates slowly enough around each test site to lead to non-vanishing bias. Inconsistency of both 1NN and the holdout can be seen as corollaries of Proposition 4.4; for 1NN, choose: $\forall n,k_{n}=1$. For the holdout, choose: $\forall n,k_{n}=n$.
@@ 5 A Consistent Estimator
We next provide a novel bound on the test risk estimation error of kNN. We propose using a kNN estimator with $k$ chosen by optimizing our bound. We show that our proposed estimator is consistent for test risk under infill asymptotics. We here focus on error estimation; in Appendix C we provide promising results for model selection and discuss open challenges.
@@ 5.1 Our Bound and Estimator
Consider a validation set $D^{\textup{{val}}}$ of size $N^{\textup{{val}}}$ and a test set $D^{\textup{{test}}}$ of size $M^{\textup{{test}}}$. Take the $k$-nearest neighbors test-risk estimator from Definition 4.2. Choose $\delta\in(0,1)$ and $k$ such that $1\leq k\leq N^{\textup{{val}}}$. Let $\rho_{k}\coloneqq\zeta^{k}(S^{\textup{{val}}}_{1:N^{\textup{{val}}}},S^{\textup{{test}}}_{1:M^{\textup{{test}}}})$ and $\beta_{\delta}:=\Delta\sqrt{\frac{1}{2}\log\frac{2}{\delta}}$. Take Assumptions 2.2, 2.1, 2.4 and 2.5. Then, with probability at least $1-\delta$,
$\smash{|R_{Q^{\textup{{test}}}}(h)-\hat{R}_{\textup{{NN}},k}(h)|}\,\,$ $\smash{\leq L\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}}\leq L\rho_{k}+\beta_{\delta}\sqrt{\max_{1\leq n\leq N^{\textup{{val}}}}Q^{\textup{{test}}}(B(S^{\textup{{val}}}_{n},\rho_{k}))/k},$ (3)
@@
where $B(S,r)$ denotes the ball of radius $r$ centered at $S$. See Assumptions 2.1, 2.5 and 2.3 for $\Delta$, $L$, $Q^{\textup{{test}}}$ respectively.
@@
We prove Theorem 5.1 in Section D.4. We use the right-hand side of the first inequality algorithmically, and the right-hand side of the second inequality to gain intuition for cases under which the bound is small, as well as in proofs. The first term on the far-right-hand side in Equation 3 is a worst-case upper bound on the bias of our estimator; it is large if the average loss varies quickly in space or if validation data is not available near test data. Larger $k$ may increase the first term. The second term comes from applying a tail bound; if most of the weight is put on a few validation points, the resulting estimator has high variance and this term is large. Sufficiently large $k$ will decrease the second term. If we can find a $k$ such that both (a) the distance from each test point to its $k$ nearest neighbors is small and (b) no validation point has too much impact on our estimator, $\smash{\hat{R}_{\textup{{NN}},k}(h)}$ provides a good estimate for $\smash{R_{Q^{\textup{{test}}}}(h)}$.
@@
Theorem 5.1 is closely related to Portier et al. (2023, Prop. 4). While there are technical differences in the proof and algorithm (and the risk that is bounded), the substantive distinction is that we state our bound directly in terms of the fill distance, instead of upper bounding this distance again as done in Portier et al. (2023). We can therefore avoid making assumptions about the distributions of the sites; we instead highlight the fill distance of the validation set as an essential quantity in controlling the accuracy of nearest neighbor risk estimation.
@@
Selection Procedure with Unknown Lipschitz Constant. If the Lipschitz constant of the average loss, $L$, can be upper bounded, for example by knowledge about how quickly varying the spatial processes involved in the analysis are, then $k$ can be selected by minimizing the first upper bound in Equation 3. Since the bound is conditional on the validation and test sites, the bound still holds with the same probability for $k$ selected by this minimization. However, it will generally be the case that the Lipschitz constant is unknown. We therefore suggest choosing the number of neighbors by minimizing the upper bound from Theorem 5.1 with $1$ in place of the Lipschitz constant:
$\smash{k^{\star}_{T}\in\arg\min\nolimits_{k\in T}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}}.$ (4)
@@
For computational efficiency, we focus on choosing $k$ as a power of 2: $T=T_{2}:=\{2^{i}\}_{i=1}^{\smash{\lfloor\log_{2}N^{\textup{{val}}}\rfloor}}$. We call the resulting estimator spatial nearest neighbors (SNN).
@@ 5.2 Our Method is Consistent
Let $\mathcal{S}=[0,1]^{d}$. Take Assumptions 2.2, 2.1, 2.4 and 2.5. Let $\tilde{\rho}:=\zeta(S^{\textup{{val}}}_{1:N},\mathcal{S})$. Let $k^{\star}_{T_{2}}\in\arg\min_{k\in T_{2}}\rho_{k}+\beta_{\delta}\|w^{\textup{{NN}},k}\|_{2}$ with $\delta=\min(1,r)$ and $r\in[\smash{c\tilde{\rho},C\tilde{\rho}}]$ for some constants (possibly depending on dimension) $c,C>0$. Then the $k^{\star}_{T_{2}}$-nearest neighbor risk estimator is consistent under infill asymptotics.
See Section D.5 for a proof of Corollary 5.2. Corollary 5.2 states that selecting the number of neighbors by minimizing an upper bound on our error in estimation leads to an estimator that is consistent regardless of the test data, as long as the validation data are dense on the unit cube. In Section D.5.4, we provide a computationally efficient algorithm for calculating an $r$ satisfying the condition $c\tilde{\rho}\leq r\leq C\tilde{\rho}$, and we prove the correctness of this algorithm.